# Rethinking Feature Backbone Fine-tuning for Remote Sensing Object Detection

### v1

@@ III Experiments
In this session, we show our novel feature backbone fine-tuning method, DBF enables more accurate model learning while significantly reducing computational costs. Please refer to the supplementary material for further information.
III-A Experimental setup
@@
Baseline architectures. Deep learning models for object detection are conventionally split into one-stage and two-stage methods. Besides, anchor-based and anchor-free methods have been recently increasingly highlighted in the classification of detector models. In this study, we opted for Faster R-CNN [24], RetinaNet [25], and FCOS [26] as the representative models in anchor-based two-stage, anchor-based one-stage, and anchor-free one-stage object detection. For all benchmark models, we considered two exemplary neural networks as feature backbones: ResNet-50 [21] and Swin-S [22], one of the celebrated convolutional neural networks and transformer architectures, respectively. We leverage the ImageNet pre-trained weights for remote sensing object detection as they are still competitive baseline models to be readily online [8]. Moreover, FPN (Feature Pyramid Network)-1x [27] is used as a neck to enhance multi-scale invariance for our models.
@@
Training details. We chose PyTorch and MMDetection / MMRotate [28] for implementing our method and baseline architectures. For model optimization, SGD (Stochastic Gradient Descent) is used with momentum of weight 0.9, weight decay 0.0001, and batch size 8. The initial learning rate is set to 0.005 for all detectors. To enhance the stability and convergence of our training process, we utilize a linear learning rate warm-up strategy over the initial 500 iterations, during which the learning rate increases linearly from zero to 33% of the maximum learning rate. The learning rate decayed by 1/4 after the 12th epoch. To prevent gradient explosion, L2 norm-based gradient clipping is used with a maximum norm of 35. We apply the following data transformations for training: ‘RResize’, ‘RRandomFlip’, ‘Normalize’, and ‘Pad’. For testing, ‘MultiScaleFlipAug’ is used as well. For a fair comparison, we set the value of the random seed to 0.
@@
Table I indicates the experimental configuration parameters in this work. As mentioned in Sec. II-C, we use Step Freezing Scheduler for DBF. Besides, we consider four $\rho$ values, $\{2,5,10,\infty\}$ in ours: the models fully fine-tuned (i.e. $\rho=1$) for the first 50 or 400 epochs are fine-tuned again with DBF.
TABLE II: Experimental results (Bold text indicates the best while underlined text means the second-best in this paper).
(c) Ours (DBF)
(a) Full training (Baseline) (b) Frozen Backbone [19] ($\rho=1\rightarrow$) $\rho=2$ ($\rho=1\rightarrow$) $\rho=5$ ($\rho=1\rightarrow$) $\rho=10$ ($\rho=1\rightarrow$) $\rho=\infty$
Dataset Detector Backbone mAP mAP $\Delta$TFLOPs mAP $\Delta$TFLOPs mAP $\Delta$TFLOPs mAP $\Delta$TFLOPs mAP $\Delta$TFLOPs
ResNet-50 [21] 69.00 61.31 (-7.69) -3,077,501 69.03 (+0.03) -1,170,338 69.22 (+0.22) -1,872,541 69.27 (+0.27) -2,106,608 69.28 (+0.28) -2,340,676
Faster R-CNN [24] Swin-S [22] 71.23 68.14 (-3.09) -6,115,002 71.80 (+0.57) -1,538,750 71.73 (+0.50) -2,462,001 71.74 (+0.51) -2,769,751 71.73 (+0.50) -3,077,501
ResNet-50 [21] 64.19 57.29 (-6.90) -3,077,501 64.48 (+0.29) -1,170,338 64.89 (+0.70) -1,872,541 64.77 (+0.58) -2,106,608 64.98 (+0.79) -2,340,676
RetinaNet [25] Swin-S [22] 64.89 61.61 (-3.28) -6,115,002 64.92 (+0.03) -1,538,750 65.02 (+0.13) -2,462,001 65.01 (+0.12) -2,769,751 64.97 (+0.08) -3,077,501
ResNet-50 [21] 68.41 59.82 (-8.59) -3,077,501 68.41 (+0.00) -1,170,338 68.80 (+0.39) -1,872,541 68.98 (+0.57) -2,106,608 69.18 (+0.77) -2,340,676
DOTA [14] FCOS [26] Swin-S [22] 71.47 67.45 (-4.02) -6,115,002 71.46 (-0.01) -1,538,750 71.44 (-0.03) -2,462,001 71.62 (+0.15) -2,769,751 71.51 (+0.04) -3,077,501
ResNet-50 [21] 63.82 54.17 (-9.65) -491,825 64.13 (+0.31) -184,434 63.85 (+0.03) -295,095 63.86 (+0.04) -331,982 63.96 (+0.14) -368,869
Faster R-CNN [24] Swin-S [22] 67.68 62.64 (-5.04) -4,588,368 67.97 (+0.29) -1,720,638 67.99 (+0.31) -2,753,021 68.06 (+0.38) -3,097,149 68.03 (+0.35) -3,441,276
ResNet-50 [21] 59.09 53.88 (-5.21) -491,825 59.26 (+0.17) -184,434 59.25 (+0.16) -295,095 58.94 (-0.15) -331,982 58.97 (-0.12) -368,869
RetinaNet [25] Swin-S [22] 65.11 60.81 (-4.30) -4,588,368 65.15 (+0.04) -1,720,638 65.14 (+0.03) -2,753,021 64.86 (-0.25) -3,097,149 64.75 (-0.36) -3,441,276
ResNet-50 [21] 61.23 55.84 (-5.39) -491,825 61.39 (+0.16) -184,434 61.63 (+0.40) -295,095 61.66 (+0.43) -331,982 61.85 (+0.62) -368,869
DIOR-R [15] FCOS [26] Swin-S [22] 67.58 67.48 (-0.10) -4,588,368 67.69 (+0.11) -1,720,638 68.08 (+0.50) -2,753,021 68.05 (+0.47) -3,097,149 68.05 (+0.47) -3,441,276
Figure 4: Training loss and validation mAP curves on DOTA and DIOR-R (Best viewed in color and zoom in).
III-B Evaluation metrics
mAP (mean Average Precision) is chosen for verifying both the precision and recall of the models across different categories in multiple object detection. Precisely, ‘mAP@50’ [29] is used for model evaluation following [30].
FLOPs (FLoating point OPerations) is adopted to measure the efficiency of the training scheme on various hardware platforms by counting the number of unit operations in each network layer. Typically, most papers calculate FLOPs based solely on the one forward pass with a batch size of 1.
On the other side, we report ${\text{F}_{\text{total}}}$ instead as follows:
@@ III-B Evaluation metrics
III-C Datasets
@@
We perform experiments on following datasets for generic multiple object detection in remote sensing: DOTA [14] and DIOR-R [15]. For our experiments, we only consider labels for oriented object detection, as detecting rotated objects is more practical in remote sensing. DOTA consists of 15 object categories including ‘car’, ‘ship’, and ‘harbor’, with 2,806 aerial images and 188,282 bounding box labels. DIOR-R surpasses DOTA with its 20 object categories and 192,512 instances captured in 23,463 images. Following [30], we showcase experimental results on the validation split.
III-D Experimental results
As seen in Table II, our method performs better in most cases for all datasets. Here, $\Delta$TFLOPs denotes the difference in Tera FLOPs relative to the baseline (a). In other words, the lower $\Delta$TFLOPs, the better. As mentioned before, the FLOPs should be calculated as ${\text{F}}_{\text{total}}$ in Eq. 1. Besides, ‘($\rho=1\rightarrow$) $\rho=\alpha$’ in (c) means that the backbone is fine-tuned with $\rho=1$ for the first 50 or 400 epochs and then fine-tuned again with $\rho=\alpha$ according to Table I. Although (a) permits adjusting the backbone for every epoch, it mostly results in slightly lower mAP than ours. Besides, (b) decreases $\Delta$TFLOPs the most, but it undergoes a significant drop in mAP (about 3 to 10%p). Compared to (a) and (b), the proposed approach balances low time complexity and precise inference ability: ours achieves the highest mAP for almost all cases, while effectively reducing computational costs during training.
@@ III-D Experimental results
To discuss the practical effectiveness of the proposed technique, we present how much training time can be saved when adopting our approach. Let the detector be FCOS with ResNet-50 as a backbone. For the training split of DOTA, using the Nvidia RTX A5000 (GPU) it takes an average of 23 minutes to fully fine-tune the model, including the backbone, in one epoch. Freezing the backbone takes 16 minutes on average. Based on these, we can approximately calculate the overall training time for each training policy as shown in Table III.
TABLE III: Examples of training time savings on Nvidia RTX A5000.
FCOS [26] + ResNet-50 [21] on DOTA [14]
@@
mAP $\Delta$TFLOPs Training time (T) $\Delta$T
$\sim$50 epochs $\sim$400 epochs
(a) Full training
$\rho=1$ $\rho=1$ 68.41 - 9,200m -
(b) Frozen Backbone [19]
$\rho=\infty$ $\rho=\infty$ 59.82 -3,077,501 6,400m -2,800m
$\rho=1$ $\rho=2$ 68.41 -1,170,338 7,975m -1,225m
$\rho=1$ $\rho=5$ 68.80 -1,872,541 7,240m -1,960m
$\rho=1$ $\rho=10$ 68.98 -2,106,608 6,995m -2,205m
(c) Ours (DBF) $\rho=1$ $\rho=\infty$ 69.18 -2,340,676 6,750m -2,450m
Here, $\Delta$T means the difference in training time compared to (a). Interestingly, ours with $\rho=\infty$ achieves a 0.77 higher mAP while reducing the training time by 2,450 minutes (1.7 days) compared to the baseline. Reducing model learning time by several hours or days holds significant benefits as follows:
  • Productivity improvement: Shorter training times enable more experiments (e.g. parameter tuning) to be conducted quickly;
  • Sustainable AI: Lower power consumption translates to reduced carbon emissions, leading to environmental conservation.
@@
Fig. 4 further demonstrates the effectiveness of the proposed method. In all subfigures, a downward trend is visible in the validation mAP curves for (a) full training, while the training loss also decreases the fastest, indicating over-fitting. On the other hand, the proposed method gradually reduces the training loss while improving the validation mAP, or at least preventing it from decreasing, especially for higher $\rho$. These observations show that the proposed method not only reduces memory usage but also helps in stable model convergence.
III-E Parameter study of $\rho$
This subsection examines how users can effectively select $\rho$ value. In the experiments, we consider four $\rho$ values, $\{2,5,10,\infty\}$ for our method. One certain thing is that as the $\rho$ value increases, more time and memory are saved in training. However, larger $\rho$ deprives the learning opportunity of the feature backbone, which might hurt the model generalization. To find the best $\rho$, we compute the average of $\Delta$mAP over all the experiments for each $\rho$. As demonstrated in Table. IV, it appears that $\rho=\infty$ is the optimal choice for better prediction.
TABLE IV: Average of $\Delta$mAP for each $\rho$ in DBF.
$\Delta$mAP $\rho=2$ $\rho=5$ $\rho=10$ $\rho=\infty$
mean ± standard deviation 0.17±0.2 0.28±0.2 0.26±0.3 0.30±0.3
IV Conclusions and Future Work

### v4

@@ 4 Experiments
In this section, we show our novel feature backbone fine-tuning method, DBF enables more accurate model learning while significantly reducing computational costs under long-term training policies with CNN-based architectures for oriented object detection in remote sensing. Note that CNN-based detectors can embrace the transformer as a feature backbone.
4.1 Experimental setup
@@
Baseline architectures. Deep learning models for object detection are conventionally split into one-stage and two-stage methods. Besides, anchor-based and anchor-free methods have been recently increasingly highlighted in the classification of detector models. In this study, we opted for Faster R-CNN [36], RetinaNet [52], and FCOS [33] as the representative models in anchor-based two-stage, anchor-based one-stage, and anchor-free one-stage object detection. For all benchmark models, we considered two exemplary neural networks as feature backbones: ResNet-50 [49] and Swin-S [50], one of the celebrated convolutional neural networks and transformer architectures, respectively. Moreover, FPN (Feature Pyramid Network)-1x [53] is used as a neck to enhance multi-scale invariance for our models. Here, we leverage ImageNet pre-trained weights for remote sensing object detection as they are still competitive baseline models to be readily online [21]. While alternative pre-trained models designed for remote sensing applications [19, 20] are available, they lack the ability to promptly adapt to novel network architectures. That is, for newly proposed neural networks, the weights trained on natural image datasets are more easily accessible. In contrast, those specialized in remote sensing imagery take considerable time to deploy. Thus, our experiments explore when the ImageNet pre-trained weights meet our method to boost remote sensing object detection.
@@
Training details. We chose PyTorch and MMDetection [54] / MMRotate [55] for implementing our method and baseline architectures. For model optimization, SGD (Stochastic Gradient Descent) [56, 57] is used with momentum of weight 0.9, weight decay 0.0001, and batch size 8. The initial learning rate is set to 0.005 for all detectors. To enhance the stability and convergence of our training process, we utilize a linear learning rate warm-up strategy [58] over the initial 500 iterations, during which the learning rate increases linearly from zero to 33% of the maximum learning rate. The learning rate decayed by 1/4 after the 12th epoch. To prevent gradient explosion, L2 norm-based gradient clipping [59] is used with a maximum norm of 35. We apply the following data transformations for training: ‘ [^1] RResize is a rotation-aware resizing operation designed for resizing images and their associated oriented bounding boxes. For DOTA, it adapts the input to a predefined size, 1024x1024 while maintaining the aspect ratio. Similarly, for DIOR-R, it adjusts all inputs to a uniform target size, 800x800 while maintaining the aspect ratio. RResize is critical for efficient batching and optimized memory usage.RResize’, ‘ [^2] RRandomFlip is a rotation-aware data augmentation technique that randomly flips images along specified axes (horizontal, vertical, or diagonal directions) while ensuring that the corresponding oriented bounding boxes are updated accordingly. For both DOTA and DIOR datasets, we consider all three directions with probability 0.25 in RRandomFlip.RRandomFlip’, and ‘ [^3] All images are normalized per each color channel (R/G/B) with means [123.68, 116.28, 103.53] and standard deviations [58.40, 57.12, 57.38].Normalize’ [55]. For a fair comparison, we set the value of the random seed to 0 during experiments.
@@
Table 1 indicates the experimental configuration parameters in this work. As mentioned in Sec. 3.3, we use Step Freezing Scheduler for DBF. Besides, we consider four $\rho$ values, $\{2,5,10,\infty\}$ in our method: the models fully fine-tuned (i.e. $\rho=1$) for the first 50 or 400 epochs are fine-tuned again with DBF under the four $\rho$ values.
Table 2: Experimental results on DOTA and DIOR-R with various convolutional object detection architectures (Bold text indicates the best while underlined text means the second-best in this paper). Notably, our method achieves higher validation mAP scores across nearly all cases for both datasets compared to approaches (a) and (b), underscoring its efficacy in long-term training scenarios for CNN-based object detection models in remote sensing.
[…]
(a) Full training
(b) Frozen Backbone
[31]
@@
ResNet-50
[49]
69.00
61.31
(-7.69)
-3,077,501
69.03
(+0.03)
-1,170,338
69.22
(+0.22)
-1,872,541
69.27
(+0.27)
-2,106,608
69.28
(+0.28)
-2,340,676
Faster R-CNN
[36]
Swin-S
[50]
71.23
68.14
(-3.09)
-6,115,002
71.80
(+0.57)
-1,538,750
71.73
(+0.50)
-2,462,001
71.74
(+0.51)
-2,769,751
71.73
(+0.50)
-3,077,501
ResNet-50
[49]
64.19
57.29
(-6.90)
-3,077,501
64.48
(+0.29)
-1,170,338
64.89
(+0.70)
-1,872,541
64.77
(+0.58)
-2,106,608
64.98
(+0.79)
-2,340,676
RetinaNet
[52]
Swin-S
[50]
64.89
61.61
(-3.28)
-6,115,002
64.92
(+0.03)
-1,538,750
65.02
(+0.13)
-2,462,001
65.01
(+0.12)
-2,769,751
64.97
(+0.08)
-3,077,501
ResNet-50
[49]
68.41
59.82
(-8.59)
-3,077,501
68.41
(+0.00)
-1,170,338
68.80
(+0.39)
-1,872,541
68.98
(+0.57)
-2,106,608
69.18
(+0.77)
-2,340,676
DOTA
[13]
FCOS
[33]
Swin-S
[50]
71.47
67.45
(-4.02)
-6,115,002
71.46
(-0.01)
-1,538,750
71.44
(-0.03)
-2,462,001
71.62
(+0.15)
-2,769,751
71.51
(+0.04)
-3,077,501
ResNet-50
[49]
63.82
54.17
(-9.65)
-491,825
64.13
(+0.31)
-184,434
63.85
(+0.03)
-295,095
63.86
(+0.04)
-331,982
63.96
(+0.14)
-368,869
Faster R-CNN
[36]
Swin-S
[50]
67.68
62.64
(-5.04)
-4,588,368
67.97
(+0.29)
-1,720,638
67.99
(+0.31)
-2,753,021
68.06
(+0.38)
-3,097,149
68.03
(+0.35)
-3,441,276
ResNet-50
[49]
59.09
53.88
(-5.21)
-491,825
59.26
(+0.17)
-184,434
59.25
(+0.16)
-295,095
58.94
(-0.15)
-331,982
58.97
(-0.12)
-368,869
RetinaNet
[52]
Swin-S
[50]
65.11
60.81
(-4.30)
-4,588,368
65.15
(+0.04)
-1,720,638
65.14
(+0.03)
-2,753,021
64.86
(-0.25)
-3,097,149
64.75
(-0.36)
-3,441,276
ResNet-50
[49]
61.23
55.84
(-5.39)
-491,825
61.39
(+0.16)
-184,434
61.63
(+0.40)
-295,095
61.66
(+0.43)
-331,982
61.85
(+0.62)
-368,869
DIOR-R
[14]
FCOS
[33]
Swin-S
[50]
67.58
67.48
(-0.10)
-4,588,368
67.69
(+0.11)
-1,720,638
68.08
(+0.50)
-2,753,021
68.05
(+0.47)
-3,097,149
68.05
(+0.47)
-3,441,276
Figure 4: Learning curves on DOTA and DIOR-R with Faster R-CNN and Swin-S: It can be seen that while our DBF shows slower training loss reduction than ‘Full training’, it shows stable and highest validation mAP values. Particularly, the validation mAP consistently shows an upward trend when using our method under long-term training scenarios. In the case of ‘Frozen Backbone’, while it incurs the lowest training cost, there is a considerable degradation in model performance compared to other approaches.
4.2 Evaluation metrics
mAP (mean Average Precision) is a key metric used to evaluate the accuracy of object localization and classification. We adopt mAP to provide a balanced assessment of models with both precision and recall across different categories in multiple object detection. Assume that TP means true positives, FP means false positives, and FN means false negatives. Precision is the proportion of predicted objects that are correctly identified as follows:
${\text{Precision}}=\frac{\text{TP}}{\text{TP+FP}}.$ (1)
Recall is the proportion of actual objects correctly detected by the model as:
${\text{Recall}}=\frac{\text{TP}}{\text{TP+FN}}.$ (2)
We recommend to see [60] for detailed calculation of mAP.
Precisely, ‘mAP@50’ [61] is used for model evaluation following [47]. In mAP@50, predictions with IoU $\geq$ 0.5 are considered correct, where IoU (Intersection over Union) per each prediction is calculated as follows:
[…]
where $\hat{y}$ is the predicted bounding box and $y$ is the ground truth bounding box.
FLOPs (FLoating point OPerations) refers to the total number of floating-point operations performed by a model during training or inference. It represents the computational cost of a model. Matrix multiplications and additions contribute to the FLOPs count during input propagation through the model. Typically, FLOPs are calculated based solely on the one forward pass with a batch size of 1 [62].
On the other side, we use $\text{F}_{\text{total}}$} instead as follows:
${\text{F}_{\text{total}}}=\sum_{\text{e}=1}^{\text{E}}\sum_{\text{i}=1}^{\text{N}}\sum_{\text{l}=1}^{\text{L}}\text{(}{\text{F}_{\text{forward}}}\text{(e, l)}+{\text{F}_{\text{backward}}}\text{(e, l)}\text{)},\vskip-2.84544pt$ (4)
where E is total training epochs, N is the number of training samples, and L is the number of network layers. ${\text{F}_{\text{forward}}}\text{(e, l)}$ and ${\text{F}_{\text{backward}}}\text{(e, l)}$ correspond to FLOPs for forward and backward operations at the e-th epoch in the l-th layer, respectively. During computation, $\text{F}_{\text{forward}}$} only focuses on the input, whereas $\text{F}_{\text{backward}}$} takes into account the weights and inputs for gradient flow. Hence, this work assumes that $\text{F}_{\text{backward}}$} is twice that of $\text{F}_{\text{forward}}$}. Intuitively, $\Delta{\text{F}_{\text{total}}}$ effectively shows how much computation reduces during all training epochs.
4.3 Datasets
@@
We perform experiments on the following datasets for generic multiple object detection in remote sensing: DOTA-v1.0 [13] and DIOR-R [14]. For our experiments, we only consider labels for oriented object detection, as detecting rotated objects is more practical in remote sensing. DOTA consists of 15 object categories including ‘car’, ‘ship’, and ‘harbor’, with 2,806 aerial images and 188,282 bounding box labels. DIOR-R surpasses DOTA with its 20 object categories and 192,512 instances captured in 23,463 images. Note that we adopt standard training and test protocols for both datasets following prior researches like [47].
4.4 Experimental results
[…]
@@ 4.4 Experimental results
To discuss the practical effectiveness of the proposed technique, we present how much training time can be saved when adopting our approach. Let the detector be FCOS with ResNet-50 as a backbone. For the training split of DOTA, using the Nvidia RTX A5000, it takes an average of 23 minutes to fully fine-tune the model, including the backbone, in one epoch. Freezing the backbone takes 16 minutes on average. Based on these, we can approximately calculate the training time for each training policy as shown in Table 3.
[…]
  • Productivity improvement: Shorter training times enable more experiments (e.g. hyper-parameter tuning or ablation study) to be conducted quickly;
  • Sustainable AI: Lower power consumption for model training translates to reduced carbon emissions, leading to environmental conservation and eco-friendly AI.
Table 3: Examples of training time savings on Nvidia RTX A5000. Although long-term training usually entails significant computational time and memory usage, the proposed approach enables a more efficient process with reduced resource requirements. Remarkably, it achieves this while delivering superior inference accuracy.
@@
FCOS +
ResNet-50 on DOTA
$\sim$50
epochs
$\sim$400
epochs
mAP $\Delta$ TFLOPs
Training Time
$\Delta$ T
(a) Full training $\rho=1$ $\rho=1$ 68.41 - 9,200m -
(b) Frozen Backbone $\rho=\infty$ $\rho=\infty$ 59.82 -3,077,501 6,400m -2,800m
[…]
@@
Fig. 4 further demonstrates the effectiveness of the proposed method. In all subfigures, a downward trend is visible in the validation mAP curves for (a) full training, while the training loss also decreases the fastest, indicating over-fitting. On the other hand, the proposed method gradually reduces the training loss while improving the validation mAP, or at least preventing it from decreasing, especially for higher $\rho$. These observations show that the proposed method helps in stable model convergence while reducing memory usage.
4.5 Parameter study of $\rho$
[…]
mean ± standard deviation 0.17±0.2 0.28±0.2 0.27±0.3 0.30±0.3
4.6 Weakness analysis
Figure 5: Effects of further training of the proposed method for RetinaNet on DIOR-R: It can be confirmed that though our DBF achieves worse mAP than the baseline when $\rho\in\{10,\infty\}$ during the initial training following the policies in Table 1, the prediction accuracy of ours are eventually improved with longer additional training. Besides, it brings steady performance improvement even in long-term training, compared to other methods.
Although our approach usually achieves competitive performance on several benchmarks and detectors, a relatively larger performance drop (-0.12 to -0.36) can be observed on the DIOR-R dataset when $\rho$ is greater than 10 for the RetinaNet architecture as shown in Table 1. Fortunately, this is not a limitation of the proposed method but because we terminate training early following the pre-fixed policy in Table 1, even though the model has not converged sufficiently. As shown in Fig 5, our method surpasses the other approaches for all $\rho$ values when we additionally train the models with ResNet-50 and Swin-S for 100 and 200 epochs more, respectively. Remark that we attempted to train more than 100 or 200 epochs for each case; but, we observed that the performance of all methods either deteriorated or showed no improvement, so we ceased further training.
Despite the success of our method, current work has notable weaknesses. First, the optimal $\rho$ value is not always fixed. As indicated in Table 2, $\rho=\infty$ is not the optimal for some cases. For instance, in Faster R-CNN with ResNet-50, the mean Average Precision gradually increases as $\rho$ goes to $\infty$ for DOTA (+0.03 $\rightarrow$ +0.22 $\rightarrow$ +0.27 $\rightarrow$ 0.28), whereas for DIOR-R, a U-shaped mAP trend is observed (+0.31 $\rightarrow$ +0.03 $\rightarrow$ +0.04 $\rightarrow$ +0.14). To address this, future work would include exploring more diverse freezing schedulers that automatically decide whether to update the backbone based on input samples or learning curves online.
Besides, there is still room for further improvement in our method, in terms of training cost. For example, Adapters [9] for NLP can be leveraged in our method to boost more efficient fine-tuning without updating the entire backbone.
5 Comparison with state-of-the-art methods
In Section 4, DBF has proven to be an efficient fine-tuning strategy for long-term training in CNN-based remote sensing object detection models. To further validate its effectiveness, we extend our experiments by incorporating DBF into a state-of-the-art end-to-end transformer-based object detection model, OrientedFormer [3]. In this section, we also include a comparison with state-of-the-art methods, for comparative evaluation.
Table 5: Comparison with state-of-the-art oriented object detection models on DIOR-R. Here, OrientedFormer, combined with our training scheme, achieves the best mAP, demonstrating the effectiveness of the proposed method.
Detector Backbone mAP
One-stage:
RetinaNet (ICCV 2017) [52] ResNet-50 [49] 57.55
DFNet (IEEE TGRS 2024) [63] ResNet-50 [49] 62.11
Oriented Rep (CVPR 2022) [34] ResNet-50 [49] 66.71
DCFL (CVPR 2023) [64] ResNet-50 [49] 66.80
Two-stage:
Gliding Vertex (IEEE TPAMI 2020) [65] ResNet-50 [49] 60.06
RoI Transformer (CVPR 2019) [66] ResNet-50 [49] 63.87
QPDet (IEEE TGRS 2023) [67] ResNet-50 [49] 64.20
AOPG (IEEE TGRS 2022) [14] ResNet-50 [49] 64.41
DODet (IEEE TGRS 2022) [68] ResNet-50 [49] 65.10
End-to-end:
ARS-DETR (IEEE TGRS 2024) [69] ResNet-50 [49] 66.12
OrientedFormer (IEEE TGRS 2024) [3] Swin-S [50] 68.84
[…]
5.1 Experimental setup
In this section, we choose OrientedFormer [3] as the representative end-to-end transformer-based object detection models in remote sensing. In addition, we opt for a DIOR-R dataset, one of the well-known large-scale datasets for oriented object detection in remote sensing. Following [3], we use the same data augmentation strategies for DIOR-R, as mentioned in Section 4. To show the accuracy of our method, we use ‘mAP@50’ as an evaluation metric, as in Section 4.2. In the original paper [3], the model is trained for 12 epochs with the image size 800x800 with DIOR-R. However, we train the model for 250 epochs with the proposed fine-tuning approach. In the previous section, we observed ‘($\rho=1\rightarrow$) $\rho=\infty$’ in DBF is the optimal choice under long-term training. Hence, we set $\rho$ as 1 for first 50 epochs, and then $\infty$ for the remaining 200 epochs.
5.2 Experimental results
We compare our approach against both recent CNN-based and transformer-based detectors on DIOR-R. As indicated in Table 5, our approach achieves 69.38%, surpassing all compared CNN-based one-stage and two-stage detectors as well as end-to-end transformer-based detectors. Specifically, our method improves upon the baseline OrientedFormer model by +0.54 mAP, demonstrating the effectiveness of integrating Dynamic Backbone Freezing (DBF) for enhanced performance in remote sensing object detection tasks.
6 Conclusions and future work